How do I solve this system of equations: 4x+3y=10
5x-y=3
by using matrices and backsolving?

Answer by scott8148(5892)

easiest way is elimination __ multiply second equation by 3 and add to first

4x+3y=10
15x-3y=9
________ __ this ELIMINATES the y
19x=19

so x=1

4(1)+3y=10 __ 3y=6 __ y=2